Circular No. 7086

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999B IN UGC 7189 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 16.9) on an unfiltered CCD frame taken on Jan. 14.29 UT utilizing the 0.35-m Tenagra I automated supernova patrol telescope in Sonoita, AZ (internet-controlled from Cottage Grove). SN 1999B is located at α = 12^{h}11^{m}16^{s}.11, δ = +74^{o}48′06′′.9 (equinox 2000.0), which is ∼ 2′′.4 west and 20′′.1 south of the nucleus of UGC 7189. The new object was confirmed on a frame taken on Jan. 15.17. No star is present at this position on frames taken on 1998 July 6.34 and Aug. 2.29 using the same equipment (limiting mag ∼ 19). Schwartz adds that the object does not appear on either the first or the second Palomar Sky Survey scans, and that it is absent from the USNO A1.0 catalogue.

SUPERNOVA 1999A IN NGC 5874 A. V. Filippenko, University of California at Berkeley, reports that

inspection of a relatively noisy CCD spectrum of SN 1999A (cf. IAUC 7083), obtained on Jan. 12 by F. Hamann (University of California, San Diego) with the 3-m Shane reflector at Lick Observatory, reveals that the object is probably a type-II supernova, with some indication of strong absorption lines of s-process elements (Ba, Sc), as was the case for SN 1987A. Moreover, the object appears to be subluminous.

V1333 AQUILAE P. J. Callanan, University College, Cork; A. V. Filippenko, University

of California at Berkeley; and M. R. Garcia, Harvard-Smithsonian Center for Astrophysics, write: “We imaged the transient low-mass x-ray binary Aql X-1 = V1333 Aql with the Near Infrared Camera on Keck I on 1998 Sept. 29.30 UT, while Aql X-1 was in quiescence according to XTE ASM observations. Our K-band images clearly resolve V1333 Aql into two stars, with a separation of 0′′.46. The stars lie approximately along the east-west direction, with the easterly star contributing 60 percent of the combined flux at K. Z-band (∼ 1.05-μm) observations indicate that this star is also somewhat bluer than its neighbor, and we speculate that this may be the true infrared/optical counterpart of Aql X-1. Further analysis is in progress to constrain the contamination of the V1333 Aql optical flux due to the line-of-sight interloper.”

1999 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green